karen left her house heading east and her sister left two hours later heading west at a rate of 5mph faster than Karen. Six hours after Karen left, she and her sister were 600 miles apart. How fast was Karen moving?
Answer by mananth(16949)

Karen speed = x
time she traveled = 6 hours
so she traveled 6x miles
Sister speed = x+5 mph
time she traveled = 4 hours
so she traveled 4(x+5) miles
4(x+5)+6x=600
4x+20+6x=600
10x=580
x=580/10
x= 58 mph
58*6=348
63*4=252
Total = 600
m.ananth@hotmail.ca